OpenStudy (anonymous):

the perimeter of a rectangular concrete slab is 114 feet and its area is 702 square feet. Find the dimensions of the rectangle

OpenStudy (anonymous):



OpenStudy (anonymous):

you have to substitude for l or w and solve the equation

OpenStudy (anonymous):

I am sorry can you break it down for me.

OpenStudy (anonymous):

lets solve for w = 702 / l the new equation is 2* (l + 702 / l ) = 114

OpenStudy (anonymous):

divide by 2 L + 702/ L = 57

OpenStudy (anonymous):

get common denominators and (L^2 + 702) / L = 57

OpenStudy (anonymous):

multiply by L both side L^2 +702 = 57 L

OpenStudy (anonymous):

now just solve quadratic equation

OpenStudy (anonymous):

Did you figure it out?

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

i got 18 and 39

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

your width is 18 and your lenght is 39

OpenStudy (anonymous):

no problem